Use of any computer software is not allowed. Solve the following problems using calculator and probability tables. Show all computations3. The tensile strength of Portland cement is being studied. Four different mixing techniques can be used economically. A completely randomized experiment was conducted and the following data were collected. Mixing Technique Tensile Strength (Ib/in) 3129 3000 2865 3200 3300 2975 2800 2900 2985 2600 2700 2600 2890 3150 3050 2765 (a) Test the hypothesis that mixing techniques affect the strength of the cement. State the model, hypotheses, test statistic, decision rule, ANOVA table and decision clearly
Answer #1
Mixing Technique yi.
1 3129 3000 2865 2890 11884
2 3200 3300 2975 3150 12625
3 2800 2900 2985 3050 11735
4 2600 2700 2600 2765 10665

H0: mu_1 = mu_2 = mu_3 = mu_4

H1: any mu_i eq 0 for i = 1, 2, 3 ,4

y.. = (11884 + 12625 + 11735 + 10665) = 46909

CF = y..2 / N = 469092/16 = 137528392.6

ssrotal-Ž Ž yリー Vij- CF = 138172041-137528392.6-643648.4

SSTreatments- ー-CF-138018132.8-137528392.6 = 489740.2

SHrror -643648.4-489740.2- 153908.2

Source of Variation SS DF MS F0
Treatments 489740.2 4-1 = 3 163246.7 12.72
Error 153908.2 12 12825.7
Total 643648.4 16-1 = 15

For alpha = 0.05,

.0.05.3, 12 3.4903

Now F0 > 3.4903, so we reject the null hypothesis and conclude with 95% confidence that atleast one mixing technique is different.

** If the answer does not match please comment.